OpenStudy (mathisfun13):

For the geometric sequence of a1=2 and r=2 find a5

OpenStudy (ash2326):

@mathisfun13 Have you studies Geometric Sequences?

OpenStudy (mathisfun13):

yea

OpenStudy (ash2326):

OK, it's of the form \[a_0, a_1, a_2...\] in terms of \(a_0\) \[a_1=a_0 \times r\] \[a_2=a_0\times r^2\] for any \(a_n\) \[a_n=a_0\times r^{n-1}\] r=common ratio Do you get this part?

OpenStudy (mathisfun13):

yesss

OpenStudy (mathisfun13):

what would the n be tho? cause i dont see a n in the problem

OpenStudy (ash2326):

For our case , we need to find a5 so n=5. Can you find a0 from a2 and given r

OpenStudy (mathisfun13):

i thinkk so

OpenStudy (ash2326):

Good, let me know if you get stuck

OpenStudy (mathisfun13):

oooh i got 32?

OpenStudy (ash2326):

Sorry, here a1 is actually a0. So it's a straight forward problem \[a_5=a_1r^{5-1}\] \[a_5=a_1r^4\] You are right :) \[a_5=2 \times 2^4=32\]